# Design and Evaluation of a PMT High-Voltage system for Deepsea Neutrino Telescope

### v1

@@ 1 Introduction
Neutrino telescopes detect high-energy neutrinos from astrophysical sources such as active galactic nuclei and core-collapse supernovae, providing probes of cosmic accelerators, multi-messenger phenomena, and fundamental physics beyond the Standard Model [Gaisser2016, Halzen:2010yj, Adrian-Martinez2016]. These instruments rely on the detection of Cherenkov photons produced by charged particles generated in neutrino interactions within transparent media such as seawater or ice. Photomultiplier tubes (PMTs) serve as the primary optical sensors, converting photon signals into electrical pulses, whose timing and charge encode the event topology and energy. Because the detectors are deployed at great depths or embedded in ice, the PMTs and their front-end electronics must operate reliably inside sealed, pressure-resistant glass vessels that preserve optical clarity and mechanical integrity. Stable and well-regulated high voltage is essential, as it directly determines the PMT gain and timing response, which in turn impacts photon hit resolution, and ultimately the quality of neutrino reconstruction. Therefore, the design of the PMT housing, its high-voltage, and readout systems plays a critical role in the overall sensitivity and long-term performance of the telescope.
@@
The TRIDENT(TRopical DEep-sea Neutrino Telescope) neutrino telescope deploys an array of hybrid Digital Optical Modules (hDOMs) anchored at depths exceeding 3,000 m [TRIDENT:2022hql]. Each hDOM integrates two complementary photon-sensing technologies—31 three-inch PMTs and 24 silicon photomultipliers (SiPMs)—to achieve high photo-coverage and directional sensitivity [trident2023hdom, Shao2025_TRIDENT]. This hybrid approach follows the multi-PMT DOM concept pioneered in KM3NeT [Kooijman2019KM3NeT]. The sensors are housed within a 17-inch-diameter borosilicate glass vessel, manufactured by Nautilus, with a 14 mm wall thickness, providing over 95 % optical transparency and sufficient mechanical strength to withstand pressures up to 670 bar [VITROVEX17]. An aluminum support frame holds all sensors in fixed orientations, while the volume between the frame and the inner wall of the glass vessel is filled with high-transparency optical gel that ensures stable optical coupling.
@@
The compact integration of multiple sensors within the 17-inch vessel imposes stringent requirements on system stability and control. Because recovery or maintenance at kilometer-scale depths is costly and technically demanding, each hDOM must operate reliably for long periods without intervention. The high-voltage (HV) system is a critical component, as its stability directly determines the PMT gain and signal-to-noise ratio, which in turn affect the detection of faint Cherenkov photons and the accuracy of event reconstruction.To meet these requirements, a CW voltage multiplier is adopted for its compact and efficient design. Its ability to generate high voltages with a minimal number of components makes it particularly suitable for the confined geometry of deep-sea optical modules  [Abbasi2010DOM].
@@
This paper presents the design and performance evaluation of a CW based high-voltage system developed for deep-sea neutrino telescope applications [CockcroftWalton1932]. The system architecture is described with an emphasis on achieving low output ripple and long-term voltage stability. Performance is validated through circuit simulations and laboratory measurements. Sec. 2 outlines the CW circuit design, control logic, and test procedures. Sec. 3 reports the PMT response, including baseline ripple and gain stability under simulated deep-sea conditions. Sec. 4 summarizes the results and discusses their relevance to future large-scale neutrino detectors.
@@ 2.1 Overall Design
The HV system adopts a distributed architecture in which each PMT is equipped with its own Cockcroft–Walton (CW) module, enabling independently adjustable bias voltages for all 31 channels in an hDOM. This one-CW-per-PMT scheme ensures uniform gain tuning, isolates channel-level variations, and avoids single-point failure modes. A dedicated control board provides channel selection and command fan-out, while the hDOM motherboard FPGA [Zhang:2025vvf] issues commands through a shared communication interface. This three-tier structure – motherboard FPGA, control board, and per-PMT CW bases – offers scalable and robust HV control. Each CW base has a compact 47 mm $\times$ 47 mm footprint, allowing all 31 modules to fit within the tight mechanical envelope of deep-sea optical modules [KM3NeT_2019DOM].
Fig. 1 shows a representative assembly of the system.The control board connects to the FPGA motherboard via a 14-pin connector and manages communication with all 31 base boards through an I2C hub. The FPGA motherboard is responsible for digitizing and processing PMT signals and communicates with the control board over the I2C interface to monitor and adjust the bias voltage. The control board distributes power and commands via the I2C hub, enabling independent channel control and simplifying the layout of the motherboard. Each base features a CW voltage multiplier circuit to generate the PMT bias voltage, an integrated ADC to monitor the output, and return voltage data to the control board for closed-loop regulation.This work focuses on the design and development of the control board and the PMT base boards.The FPGA motherboard was also designed and developed by the TRIDENT collaboration, but will not be discussed in detail here.
@@ 2.2 Base and Control Board Implementation
The control board manages voltage configuration and communication for all 31 bases through a multiplexed I2C architecture. The control board is connected to the FPGA motherboard via a 14-pin header, which also interfaces with four 8-channel I2C-bus multiplexers with a reset (PCA9547). As shown in Fig. 2, each PCA9547 is routed to eight Base connectors. Upon power-up, the PCA9547 enables Channel 0 by default. Therefore, to avoid I2C bus contention, Channel 0 on the remaining three devices is disabled before communication begins. Using the PCA9547 eliminates the need to assign 31 unique device addresses and simplifies firmware configuration. During operation, enabling one PCA9547 channel at a time selects the corresponding Base for I2C communication. Tests confirm that the FPGA motherboard can address and monitor all 31 Bases via I2C without conflict.
Voltage monitoring and control signals are transmitted back to the FPGA motherboard via the same multiplexed I2C network.After being scaled down by a precision resistor divider, the high voltage signal is sampled by an ADC.The MCU samples the high-voltage output using the on-board ADC and transmits the measurements to the control board over the I2C bus. The control board includes power-protection circuitry and four 8-channel I2C multiplexers with reset capability. Because each device powers up with Channel 0 enabled by default, the FPGA first disables all channels to avoid bus contention. A selected channel is then enabled to establish communication with the corresponding Base, allowing voltage configuration and readback through the FPGA motherboard.
Each PMT base consists of three functional blocks: an MCU, an inductive resonant driver, and a CW voltage multiplier, as shown in Fig. 2.The inductive resonant circuit is dominated by the inductors L3 and L5.The CW circuit is primarily composed of multiple diodes and capacitors.The MCU generates a pulse-width–modulated (PWM) signal whose frequency and duty cycle control the drive amplitude. Through a USB-to-serial interface, the output voltage is adjustable from 0 V to -1.5 kV, and real-time feedback from the on-board ADC provides closed-loop regulation. The resonant driver converts the DC supply into an AC waveform, while low-/high-pass filters and ferrite beads suppress switching noise[Timmer_2010][IceCube:2021_mDOM].
@@
Figure 2: Main circuit of the PMT high-voltage system
@@
The CW multiplier and resonant driver form the core of the high-voltage generation circuit on the base. The multiplier is built from cascaded diode–capacitor stages implementing a 3:1:1:1:1:1:1:1:1:1:1 dynode voltage ratio, as recommended for the Hamamatsu R14374 and NNVT N2031 PMT to reduce transition-time spread and maintain timing uniformity. A PWM waveform generated by the MCU passes through low-pass and high-pass filters to suppress switching noise before driving the LC resonant stage. Depending on the resonant parameters, the PWM frequency is set to approximately 80 kHz, producing a 100–150 Vpp AC signal that feeds the multiplier. A parallel RC filter smooths the rectified output to provide a stable DC bias for PMT operation. The bases operate inside a sealed 0.5–0.7 atm nitrogen environment within the glass sphere, where the seawater temperature remains 2–4 °C. Under these conditions, the design should produce stable high-voltage outputs with sub-millivolt baseline fluctuations.
The operation of the CW multiplier can be understood by analyzing the two phases of the input AC waveform (typically $\sim$80 kHz). During the positive half-cycle, the input polarity forward-biases a set of diodes, allowing the upper-tier capacitors (C33,C34,C35) to charge. Each capacitor charges to a voltage close to the peak value of the input AC source ($V_{\mathrm{pp}}$). During the negative half-cycle, the input polarity reverses. The previously charged upper-tier capacitors act as series voltage sources, adding to the input drive. This combined voltage forward-biases the next diode stage and charges the lower-tier capacitors (C17, C18, C19). The resulting voltage on these capacitors approaches $2V_{\mathrm{pp}}$ in the first stage, with the multiplication effect propagating through subsequent stages.
A key advantage of the CW topology is that the voltage stress on each capacitor does not exceed $2~V_{pp}$, permitting the use of capacitors with lower voltage ratings than the overall output. The high-voltage output is then routed through a filter network and series damping resistors. These resistors suppress overshoot and high-frequency ringing caused by the PMT’s output capacitance, improving pulse stability. The polarity of the output is determined by the diode orientation; for example, a positive high voltage for the PMT cathode is obtained by reversing the diode directions relative to a negative-output configuration.
@@ 2.3 Base High Voltage Test
We simulated the performance of the CW base to evaluate voltage buildup and steady-state behavior across the multiplier stages. The circuit was modeled in ADI LTspice using component values that match those of the prototype, including the resistor–capacitor network and output filter. Fig. 3 shows the simulated stage voltages as a function of time. All stages reach their steady-state values after roughly 100 ms, with uniform voltage spacing along the multiplication chain that is consistent with the design specifications.
@@
Three bases were tested to verify the dynode voltages and their consistency with the designed ratio. Each unit was placed on an insulating substrate to avoid external interference. Stage voltages from DY10 to the cathode were measured sequentially using a high-voltage probe (HVP-40, 1000:1 attenuation), with readings taken after a 5-second stabilization period. The first base yielded –104 V at DY10, which was adopted as the reference setting for the remaining units. During all measurements, the anode (P) was held at ground, and the cathode (K) reached approximately –1270 V. Table 1 lists the measured stage voltages, which follow the expected CW scaling and agree with the LTspice simulation.Under the conditions of a regulated 5 V input and a -1200 V output, the comprehensive performance of the base board was evaluated, with the key results summarized in Table 2 .
@@ 2.4 Control Logic
The PMTs operate at a typical -1250 V to achieve a gain of $10^{7}$. Because the gain follows a power-law dependence on the bias voltage, small variations in bias can lead to significant deviations in gain. The HV required for a specific PMT varies due to differences in the manufacturing processes of Hamamatsu and NNVT. To achieve a uniform response across the array, each of the 31 PMTs within an hDOM requires individually adjustable HV control. Furthermore, the gain may drift over extended periods of operation, and specific channels may need to be selectively disabled under deep-sea conditions. These considerations necessitate fine, stable, and independently configurable high-voltage control for all PMTs.
Each base MCU is programmed with a default PWM frequency (typically 80 kHz) and a 50% duty cycle, which produces an initial high-voltage output. Regulation is achieved by comparing the on-board ADC reading with the target value. Adjusting the PWM frequency provides coarse control, with each step corresponding to roughly a 100 V change in output voltage. Fig. 4 shows the relationship between the PWM parameters and the cathode voltage: the output peaks near a 50% duty cycle and decreases symmetrically toward either end in a near-parabolic profile.
After system integration, the FPGA executes a two-step control sequence. FPGA first scans the PWM drive frequency around an initial reference to reach the approximate target range, and then iteratively fine-tunes the duty cycle based on feedback from the on-board ADC until the desired high-voltage value is achieved. This closed-loop procedure enables each PMT base to autonomously converge to its operating voltage with sub-percent precision, ensuring uniform gain calibration and stable performance across all channels.
@@ 3.1 Experimental Setup
The objective of this study is to evaluate the performance and stability of PMTs equipped with soldered CW bases. Measurements focus on baseline stability, gain, and time resolution, as well as assessing environmental influences, such as temperature.
All measurements are performed in a dark room to maintain a stable optical environment. To measure the timing performance, a PMT test bench equipped with a picosecond laser is adopted, as shown in Fig. 5. A picosecond laser emitting at 532 nm served as the pulsed light source.
@@
[…]
For stability and environmental tests, in this work, we will show results from our tests on 31 PMTs in one hDOM. The 31 PMTs and the complete HV system were integrated into an hDOM and placed in a temperature-controlled chamber maintained at $2^{\circ}\mathrm{C}$. The internal volume of the hDOM was filled with nitrogen gas at a pressure of $\sim$ 0.5 atm, replicating the thermal and pressure conditions expected during deep-sea operation. Gain and baseline stability were evaluated by monitoring single-photon dark-count signals under these conditions.
3.2 PMT gain measurement and stability
The PMT gain was calibrated using single-photoelectron (SPE) charge spectra measured under low-light conditions, where the mean photoelectron occupancy satisfies $\lambda\ll 1$. The charge of each PMT pulse is obtained by integrating the waveform over a predefined time window around the pulse. The resulting charge distribution, $f(q)$, can be modeled as a Poisson-weighted sum of $n$-photoelectron contributions [DEAP:2017fgw]:

### v2

@@ 1 Introduction
Neutrino telescopes detect high-energy neutrinos from astrophysical sources such as active galactic nuclei and core-collapse supernovae, providing probes of cosmic accelerators, multi-messenger phenomena, and fundamental physics beyond the Standard Model [7, 9, 2]. Experiments like IceCube and KM3NeT primarily rely on the detection of Cherenkov photons produced by charged particles generated in neutrino interactions within transparent media such as seawater or ice. Photomultiplier tubes (PMTs) serve as the primary optical sensors, converting photon signals into electrical pulses, whose timing and charge encode the event topology and energy. Because the detectors are deployed at great depths or embedded in ice, the PMTs and their front-end electronics must operate reliably inside sealed, pressure-resistant glass vessels that preserve optical clarity and mechanical integrity. Stable and well-regulated high voltage is essential, as it directly determines the PMT gain and timing response, which in turn impacts photon hit resolution, and ultimately the quality of neutrino reconstruction. Therefore, the design of the PMT housing, its high-voltage, and readout systems plays a critical role in the overall sensitivity and long-term performance of the telescope.
@@
The TRIDENT(TRopical DEep-sea Neutrino Telescope) neutrino telescope deploys an array of hybrid Digital Optical Modules (hDOMs) anchored at depths exceeding 3,000 m [16]. Each hDOM integrates two complementary photon-sensing technologies—31 three-inch PMTs and 24 silicon photomultipliers (SiPMs)—to achieve high photo-coverage and directional sensitivity [20, 13]. This hybrid approach follows the multi-PMT DOM concept pioneered in KM3NeT [11]. The sensors are housed within a 17-inch-diameter borosilicate glass vessel, manufactured by Nautilus, with a 14 mm wall thickness, providing over 95 % optical transparency and sufficient mechanical strength to withstand pressures up to 670 bar [8]. An aluminum support frame holds all sensors in fixed orientations, while the volume between the frame and the inner wall of the glass vessel is filled with high-transparency optical gel that ensures stable optical coupling.
@@
The compact integration of multiple sensors within the 17-inch vessel imposes stringent requirements on system stability and control. Because recovery or maintenance at kilometer-scale depths is costly and technically demanding, each hDOM must operate reliably for long periods without intervention. The high-voltage (HV) system is a critical component, as its stability directly determines the PMT gain and signal-to-noise ratio, which in turn affect the detection of faint Cherenkov photons and the accuracy of event reconstruction.To meet these requirements, a CW voltage multiplier is adopted for its compact and efficient design. Its ability to generate high voltages with a minimal number of components makes it particularly suitable for the confined geometry of deep-sea optical modules  [1].
@@
This paper presents the design and performance evaluation of a CW based high-voltage system developed for deep-sea neutrino telescope applications [5]. The system architecture is described with an emphasis on achieving low output ripple and long-term voltage stability. Performance is validated through circuit simulations and laboratory measurements. Sec. 2 outlines the CW circuit design, control logic, and test procedures. Sec. 3 reports the PMT response, including baseline ripple and gain stability under simulated deep-sea conditions. Sec. 4 summarizes the results and discusses their relevance to future large-scale neutrino detectors.
@@ 2.1 Overall Design
The HV system adopts a distributed architecture in which each PMT is equipped with an independent Cockcroft–Walton (CW) base. This one-CW-per-PMT scheme ensures uniform gain tuning, isolates channel-level variations, and avoids single-point failure modes. The system follows a three-tier hierarchy: the hDOM motherboard FPGA [17] issues high-level commands; a dedicated control board manages command fan-out and channel selection; and the individual PMT bases execute voltage generation. Each base maintains a compact $47\times 47$ mm footprint, allowing all 31 PMTs to be integrated within the restricted mechanical envelope of the deep-sea optical module.
Fig. 1 shows the actual photos of the control and base boards. The FPGA motherboard digitizes PMT signals and communicates with the control board via a shared I2C interface to adjust bias voltages. The control board, acting as an intermediate hub, distributes power and commands to the bases, simplifying the motherboard’s layout. This work focuses specifically on the design and characterization of the control board and the per-PMT bases. The motherboard was also designed and developed by the TRIDENT collaboration, but it will not be discussed in detail here.
@@ 2.2 Base and Control Board Implementation
The control board manages communication between the FPGA motherboard and the 31 bases through a multiplexed $I^{2}C$ architecture. It utilizes four 8-channel $I^{2}C$ multiplexers with reset capability (PCA9547) to interface with the FPGA via a 14-pin header. As illustrated in Fig. 2, each multiplexer is routed to eight base connectors. To prevent bus contention arising from the PCA9547’s default power-up state—where Channel 0 is active—the FPGA firmware initially disables all multiplexer channels. Communication is subsequently established by enabling only the specific channel corresponding to the target base, allowing for isolated voltage configuration and data readback. This architecture eliminates the requirement for 31 unique device addresses and has been experimentally verified to provide reliable, conflict-free monitoring across all channels.
Each base comprises three functional blocks: an MCU, an inductive resonant driver, and a Cockcroft-Walton (CW) voltage multiplier (Fig. 2). The MCU generates a pulse-width–modulated (PWM) signal to drive the resonant stage, which employs inductors to produce an AC waveform. This waveform is subsequently amplified and rectified by the CW diode-capacitor network to provide an adjustable output from 0 V to 1.5-1.5 kV. Output regulation is achieved by tuning the PWM frequency and duty cycle. For closed-loop control, the high-voltage output is scaled by a precision resistor divider and digitized by the MCU’s internal ADC, providing real-time feedback for regulation.
@@
Figure 2: Schematic of the three-tier architecture for the PMT high-voltage system.
The resonant driver and CW multiplier constitute the core high-voltage generation circuit. Depending on the resonant parameters, the PWM frequency is set to approximately 80 kHz, producing a 100–150 Vpp AC signal that feeds the multiplier. The working principle of the CW multiplier can be understood by analyzing the two phases of the input AC waveform. During the positive half-cycle, the diodes are forward-biased to charge the upper-tier capacitors to the peak input voltage ($V_{\text{p}}$). Upon polarity reversal during the negative half-cycle, these capacitors act as series voltage sources that, combined with the input drive, charge the lower-tier capacitors. This results in a voltage of approximately $2V_{\text{p}}$ at the first stage, with the multiplication effect propagating through subsequent cascaded stages to achieve the required high-voltage output. A key advantage of the CW topology is that the voltage stress on individual capacitors does not exceed $2V_{\text{p}}$, permitting the use of capacitors with voltage ratings significantly lower than the total output. The polarity of the output is determined by the diode orientation; for example, a positive high voltage is obtained by reversing the diode directions relative to a negative-output configuration.
@@
The multiplier is configured to provide a 3:1:1:1:1:1:1:1:1:1:1 dynode voltage ratio. This specific distribution is recommended for the Hamamatsu R14374 and NNVT N2031 PMTs to minimize transition-time spread and maintain timing uniformity. The high-voltage output is then routed through a filter network and series damping resistors. These resistors suppress overshoot and high-frequency ringing caused by the PMT’s output capacitance, improving pulse stability. A parallel RC filter smooths the rectified output to provide a stable DC bias for PMT operation. To ensure signal integrity, low/high-pass filters and ferrite beads are integrated within the driver stage to suppress switching noise [15, 14]. The bases operate inside a sealed 0.5–0.7 atm nitrogen environment within the glass sphere, where the seawater temperature remains 2–4 °C. Under these conditions, the design should produce stable high-voltage outputs with sub-millivolt baseline fluctuations.
@@ 2.3 Base High Voltage Test
We simulated the performance of the PMT base to evaluate voltage buildup and steady-state behavior across the multiplier stages. The circuit was modeled in ADI LTspice using component values that match those of the prototype, including the resistor–capacitor network and output filter. Fig. 3 shows the simulated stage voltages as a function of time. All stages reach their steady-state values after roughly 100 ms, with uniform voltage spacing along the multiplication chain that is consistent with the design specifications.
@@
For laboratory characterization and firmware updates, a USB-to-serial interface was utilized to connect the base MCU directly to a host PC. This interface acts as a communication bridge, converting USB signals into the asynchronous serial data (UART) required by the MCU, allowing the base to be tested independently of the hDOM motherboard.
Using this setup, three bases were tested to verify dynode voltages and consistency with the designed ratio. Each unit was placed on an insulating substrate, and stage voltages from DY10 to the cathode were measured sequentially using a high-voltage probe (HVP-40, 1000:1 attenuation) after a 5-second stabilization period. The first base yielded -104 V at DY10, which was adopted as the reference setting for the remaining units. During measurements, the anode (P) was held at ground, and the cathode (K) reached approximately -1270 V. Table 1 lists the measured stage voltages, which follow the expected CW scaling and agree with the LTspice simulation. Key performance results under a regulated 5 V input and a -1200 V output are summarized in Table 2.
@@ 2.4 Control Logic
The PMTs typically operate at -1250 V to achieve a gain of $10^{7}$. Since the gain follows a power-law dependence on the bias voltage, small variations in the bias can lead to significant gain deviations. The required high voltage varies between PMTs. Furthermore, the gain may drift over extended periods of operation, and specific channels may need to be selectively disabled under deep-sea conditions. These considerations necessitate fine, stable, and independently configurable high-voltage control for all PMTs within an hDOM.
We conducted a systematic sweep of the PWM parameters using the laboratory bench setup to map the circuit’s transfer function. With the base connected to a host PC, the MCU was commanded to sweep through a range of frequencies and duty cycles. The resulting high-voltage output was sampled by the on-board ADC and transmitted back to the PC. As shown in Fig. 4, the relationship between PWM parameters and the cathode voltage exhibits resonant behavior. At a fixed frequency, the output peaks near a 50% duty cycle and follows a near-parabolic profile. Adjusting the PWM frequency provides coarse control—shifting the resonant peak to allow for approximately 100 V steps while fine-tuning the duty cycle enables precision regulation.
After system integration, the FPGA executes a two-step autonomous control sequence. The FPGA first scans the PWM drive frequency around an initial reference to reach the approximate target range, and then iteratively fine-tunes the duty cycle based on feedback from the on-board ADC until the desired high-voltage value is achieved. This closed-loop procedure enables each PMT base to autonomously converge to its operating voltage with sub-percent precision, ensuring uniform gain calibration and stable performance across all channels.
@@ 3.1 Experimental Setup
The performance of the PMTs with the bases is directly linked to the physical performance of the hDOM detector. We evaluate the stability and timing characteristics of the integrated units through two distinct experimental configurations:
  • Laser Test Bench (Individual Characterization): Used to measure parameters such as the gain-voltage curve and Transit Time Spread (TTS). As illustrated in Fig. 5, a picosecond laser ($532$ nm, $\sim 10$ ps pulse width) with an adjustable repetition rate ($500$ Hz to $300$ kHz) served as the light source. The laser output was split into a reference trigger branch, coupled to a fast photodiode, and an illumination branch. To reach the single-photoelectron (SPE) regime, the illumination passed through motorized neutral-density (ND) filter wheels (Tab. 3) and a diffuser fiber tip. Data were captured using a PicoScope 6426E ($5$ GS/s sampling rate, $1$ GHz bandwidth) triggered externally by the laser synchronization signal. This ensures that incident photons arrive at a fixed, predictable time delay relative to the trigger, facilitating high-precision timing analysis.
  • Integrated hDOM Setup (System-Level Validation): Used to assess long-term gain and baseline stability for the full array of 31 PMTs (12 NNVT, 19 Hamamatsu) integrated with the complete HV system. To simulate deep-sea conditions, the hDOM was placed in a climatic chamber at $2^{\circ}\mathrm{C}$ and backfilled with dry nitrogen to $\sim 0.5$ atm. Data acquisition in this setup supports three distinct trigger modes: (1) PMT self-trigger, which initiates a readout when a single-channel pulse height exceeds a predefined threshold; (2) periodic random trigger, used primarily for unbiased sampling of the baseline and dark count rate (DCR); and (3) local coincidence trigger (CL2), which initiates a readout when two or more PMTs exceed their thresholds within a 20 ns coincidence window. These modes allowed the FPGA motherboard to continuously monitor the integrated HV system’s stability and noise floor over a 100-hour period.
Both configurations were housed in dark rooms to maintain a stable optical environment. The PMTs were granted an adequate stabilization period to reach thermal and electronic equilibrium, ensuring the DCR had settled near the manufacturer’s reference value ($\sim$1 kHz) prior to data acquisition.
@@
3.2 PMT gain
The PMT gain was calibrated using the laser test bench to characterize the response of individual units. The procedure involves measuring single-photoelectron (SPE) charge spectra under low-light conditions, where the mean photoelectron occupancy satisfies $\lambda\ll 1$. The charge of each PMT pulse is obtained by integrating the waveform over a predefined time window around the pulse. The resulting charge distribution, $f(q)$, can be modeled as a Poisson-weighted sum of $n$-photoelectron contributions [4]: